OpenStudy (anonymous):

Sally swam a total distance of 3 and 1 over 2 miles during the months of November and December. If Sally only swam 1 over 4 of a mile every day, which expression shows the number of days she went swimming? 3 and 1 over 2 ⋅ 1 over 4 3 and 1 over 2 ÷ 1 over 4 3 and 1 over 2 + 1 over 4 3 and 1 over 2 – 1 over 4

OpenStudy (kamibug):

You want to find how many times 1/4 goes into 3.5 :) Which operation do you think you should use?

OpenStudy (michele_laino):

\[\frac{ \frac{ 7 }{ 2 } }{ \frac{ 1 }{ 4 } }\]

OpenStudy (michele_laino):

sinc 3 and 1 over 2 =\[=\frac{ 7 }{ 2 }\]

OpenStudy (michele_laino):

so?

OpenStudy (anonymous):

? im not sure @Michele_Laino

OpenStudy (michele_laino):

why?

OpenStudy (anonymous):

so 7/2...

OpenStudy (michele_laino):

please note that : if Sally swam 1 over 4 miles a day, then, she swam 2*(1/4) miles after two days, and she swan 3*(1/4) miles after days, and so on...

OpenStudy (michele_laino):

please you have multiply the first fraction, namely 7/2, times the inverse of the second fraction, namely 1/4, so you have to calculate this: \[\frac{ 7 }{ 2 }*4=...\]

OpenStudy (anonymous):

14

OpenStudy (michele_laino):

that's right! Sally swam 3 and 1/2 miles after 14 days!

OpenStudy (kamibug):

You divided 3 and 1/2 by 1/4 to get the answer. ^.^

OpenStudy (anonymous):

so then which answer would that be @Michele_Laino

OpenStudy (anonymous):

So a? Thanks @KamiBug !!!

OpenStudy (kamibug):

Actually, option A has the multiplication sign. :P It would be B.